You wish to mix wine with 25% alcohol and peach schnapps with 50% alcohol to obtain a different drink. How many liters of wine would you need to make 20 liters of the mixture with a 40% alcohol content?
Answer by funmath(2933)

You wish to mix wine with 25% alcohol and peach schnapps with 50% alcohol to obtain a different drink. How many liters of wine would you need to make 20 liters of the mixture with a 40% alcohol content?
The amount of the mixture is:20 L
Let the amount of wine be:x
Then the amount of schnapps is the rest of the mixture: 20-x
:
The amount of alcohol in the mixture is: .40(20)
The amount of alcohol in the wine is: .25x
The amount of alcohol in the schnapps is: .50(20-x)
:
You add the wine and schnapps to get the mixture, so your equation to solve is:
.25x+.50(20-x)=.40(20)
.25x+10-.50x=8
-.25x+10=8
-.25x+10-10=8-10
-.25x=-2
-.25x/-.25=-2/-.25
x=8
****(answer)The amount of wine:x=8 L
(in case you were wondering) The amount of schnapps: 20-x=20-8=12 L
Happy Calculating, and Don't Drink and Drive!!!